# Efficient Fairness-Performance Pareto Front Computation

### v1

@@ Appendix A Monotonicity of Loss Under Representations
As discussed in Section 3, we observe that representations can not increase the performance of the classifier (i.e decrease the loss).
@@
[…]
@@ Appendix B MIFPO and Optimal Transport
and this is independent of the variables $r$! Therefore we can write the cost (14) as
@@
Proof.
@@
$\begin{cases}T^{\prime}_{a}(h,s)=T_{a}(h,s)&\text{for all $a\in\left\{0,1\right\}$, all $s\in S$ and all $h\in\mathcal{Z}\setminus\left\{z\right\}$ }\\ T^{\prime}_{0}(z^{\prime},x)=T_{0}(z,x)&\\ T^{\prime}_{0}(z^{\prime\prime},u)=T_{0}(z,u)&\text{ for all $u\in U^{r}$}\\ T^{\prime}_{1}(z^{\prime},v)=\kappa T_{1}(z,v)&\text{ for all $v\in V$}\\ T^{\prime}_{1}(z^{\prime\prime},v)=(1-\kappa)T_{1}(z,v)&\text{ for all $v\in V$}.\end{cases}$ (30)
All values of $T^{\prime}$ that were not explicitly defined in (30) are set to 00. In words, on the side of $S_{0}$, we move all the parents of $z$ except $x$ to be the parents of $z^{\prime\prime}$, while $z^{\prime}$ will have a single parent, $x$. On the $S_{1}$ side, both $z^{\prime}$ and $z^{\prime\prime}$ will have the same parents as $z$, with transitions multiplied by $\kappa$ and $1-\kappa$ respectively. The multiplication by $\kappa$ is crucial for showing both inequlaities in (7).
@@
[…]
@@
$\sum_{s\in S_{1}}\mathbb{P}\left(Y=y|s,a=1\right)\mathbb{P}\left(s|a=1,z\right)=\sum_{s\in S_{1}}\mathbb{P}\left(Y=y|s,a=1\right)\mathbb{P}\left(s|a=1,z^{\prime}\right)=\sum_{s\in S_{1}}\mathbb{P}\left(Y=y|s,a=1\right)\mathbb{P}\left(s|a=1,z^{\prime\prime}\right).$ (64)
@@ Appendix D Uniform Approximation and Two Point Representations
As discussed in Sections 1,5.2, we are interested in showing that all optimal invertible representations, no matter which, and no matter on which set $\mathcal{Z}^{\prime}$, can be approximated using a representation with the following property: For every $u\in S_{0},v\in S_{1}$, there are at most $k$ points $z\in\mathcal{Z}$ that have $(u,v)$ as parents, see Figure 2(a). Here $k$ would depend only on the desired approximation degree, but not on $\mathcal{Z}^{\prime}$, or on the exact representation we are approximating. We therefore refer to this result as the Uniform Approximation result. Its implications for practical use were discussed in Section 5.2.
@@
This Section uses the notation of Section 4. Let $T$ be an invertible representation, let $u\in S_{0},v\in S_{1}$ be some points, and denote by $\mathcal{Z}_{uv}=\left\{z^{j}\right\}_{1}^{k}$ the set of all points $z\in\mathcal{Z}$ which have $u$ and $v$ as parents. Denote by
@@
the total weights of $\beta_{0}$ and $\beta_{1}$ transferred by the representation from $u$ and $v$ respectively to $\mathcal{Z}_{uv}$. Recall that $\rho_{u},\rho_{v}$ denote the distributions of $Y$ conditioned on $u,v$. We call the situation above, i.e. the collection of numbers ($\left(\left\{\beta_{0}(u)T_{0}(z^{j},u)\right\}_{j\leq k},\left\{\beta_{1}(v)T_{1}(Z^{j},v)\right\}_{j\leq k}\right)$, a two point representation, since it describes how the weight from the points $u,v$ is distributed in the representation, independently of the rest of the representation. The contribution of $\mathcal{Z}_{uv}$ to the global performance cost is
@@
For every $\varepsilon>0$, there a number $n=n_{\varepsilon}$ depending only on the function $h$, with the following property: For every two-point representation ($\left(\left\{\beta_{0}(u)T_{0}(z^{j},u)\right\}_{j\leq k},\left\{\beta_{1}(v)T_{1}(Z^{j},v)\right\}_{j\leq k}\right)$, with total weights $w_{u},w_{v}$, there is a two point representation $T^{\prime}$ on a set $\mathcal{Z}^{\prime}_{u,v}$, with the same total weights, such that $\left|\mathcal{Z}^{\prime}_{uv}\right|\leq n$, and such that
@@ Appendix E Concavity Of $E_{r}$
Appendix F Additional Proofs
F.1 Proof Of Lemma 5.1
@@ Appendix F Additional Proofs
F.2 The Factor Representation
Fix $z\in\mathcal{Z}$. To show $\mu_{0}(z)=\bar{\mu_{0}}(z)$ write, by definition,
$\bar{\mu}_{0}(z)$ $=\sum_{\rho}\bar{\beta}_{0}(\rho)\bar{T}_{0}(z,\rho)$ (106)
$=\sum_{\rho}\sum_{v\in\Psi^{-1}(\left\{\rho\right\})}\beta_{0}(v)T_{0}(z,v)$ (107)
$=\sum_{v\in S_{0}}\beta_{0}(v)T_{0}(z,v)$ (108)
$=\mu_{0}(z).$ (109)
The case $a=1$ is shown similarly. Next, the above also implies that
$\mathbb{P}_{\bar{T}}\left(z\right)=\alpha_{0}\bar{\mu_{0}}(z)+\alpha_{1}\bar{\mu_{1}}(z)=\alpha_{0}\mu_{0}(z)+\alpha_{1}\mu_{1}(z)=\mathbb{P}_{T}\left(z\right).$ (110)
Finally, by definition we have
$\mathbb{P}_{\bar{T}}\left(Y|z\right)$ $=\sum_{a\in\left\{0,1\right\}}\sum_{\rho}\mathbb{P}_{\bar{T}}\left(Y|\rho,a\right)\mathbb{P}_{\bar{T}}\left(\rho,a|z\right)$ (111)
$=\mathbb{P}_{\bar{T}}\left(z\right)^{-1}\sum_{a\in\left\{0,1\right\}}\sum_{\rho}\mathbb{P}_{\bar{T}}\left(Y|\rho,a\right)\mathbb{P}_{\bar{T}}\left(z|\rho,a\right)\bar{\beta}_{a}(\rho)\alpha_{a}$ (112)
Taking (110) into account, to show $\mathbb{P}_{\bar{T}}\left(Y|z\right)=\mathbb{P}_{T}\left(Y|z\right)$ it thus suffices to show that for a fixed $a$,
$\sum_{\rho}\mathbb{P}_{\bar{T}}\left(Y|\rho,a\right)\mathbb{P}_{\bar{T}}\left(z|\rho,a\right)\bar{\beta}_{a}(\rho)\alpha_{a}=\sum_{x}\mathbb{P}_{T}\left(Y|x,a\right)\mathbb{P}_{T}\left(z|x,a\right)\beta_{a}(\rho)\alpha_{a}.$ (113)
Let us consider $a=0$. Then
$\sum_{\rho}\mathbb{P}_{\bar{T}}\left(Y|\rho,a=0\right)\mathbb{P}_{\bar{T}}\left(z|\rho,a=0\right)\bar{\beta}_{0}(\rho)\alpha_{0}$ $=\alpha_{0}\sum_{\rho}\rho\bar{T}_{0}(z,\rho)\bar{\beta}_{0}(\rho)$ (114)
$=\alpha_{0}\sum_{\rho}\rho\sum_{v\in\Psi^{-1}(\left\{\rho\right\})}\beta_{0}(v)T_{0}(z,v)$ (115)
$=\alpha_{0}\sum_{v\in S_{0}}\rho_{v}\beta_{0}(v)T_{0}(z,v).$ (116)
where we have use the definitions on line (114), and in particular that $\mathbb{P}_{\bar{T}}\left(Y|\rho,a\right)=\rho$ by construction. This shows (113) for $a=0$, and the case $a=1$ is shown similarly, thus completing the proof.
Appendix G Experiments - Additional Details
G.1 Datasets for evaluations
The FNF data preprocessing, as implemented in Balunović et al., 2022b , was used for both MIFPO and FNF models. In running FNF, we have used the original code and configuration from Balunović et al., 2022b , with the exception that we have disabled the noise injection into features, and the subsequent logit transformation. This was disabled since this step was not a part of preprocessing, and due its particular implementation in Balunović et al., 2022b , it significantly complicated the comparison with an external model like XGBoost. This step was also performed during both training and evaluation, thus producing a less clean experiment, and was not documented anywhere in the paper Balunović et al., 2022a . We note that retaining this step would not have changed the FNF related analysis or its conclusions, as presented in Section 6.
@@ G.1 Datasets for evaluations
G.2 Minimization of Concave Functions under Convex Constraints
@@
We can write our cost function under the entropy accuracy error as:
@@
$=\mathrm{entropy}(a+b-a\cdot p_{a}-b\cdot p_{b})+\mathrm{entropy}(a\cdot p_{a}+b\cdot p_{b})-\mathrm{entropy}(a+b)$
@@
G.3 Implementations and computational details
The Pareto front evaluation requires two main parts - solving the optimization problem described above, and building a calibrated classifier required for evaluating $c_{a}=P(Y|X,A=a)$ (see Algorithm 1).
Appendix H Fair Classifiers As Fair Representations
In this Section we show that the (accuracy, statistical parity) Pareto frontier of binary classifiers can be computed from the Pareto frontier of representations with total variation fairness distance.
We then proceed to evaluate the Pareto front of a few widely used fair classifiers on standard benchmark datasets, and compare it to the Pareto front computed by MIFPO. We find that in all cases, the standard classifiers produce fronts that are significantly weaker than the MIFPO front.
H.1 The Relation between Representation and Classifier Pareto Fronts

### v2

@@ Appendix A Monotonicity of Loss Under Representations
As discussed in Section 3.1, we observe that representations can not increase the performance of the classifier (i.e decrease the loss).
@@
$\mathbb{P}\left(Y=y|Z=z\right)$ $=\sum_{a}\int dx\hskip 5.69054pt\mathbb{P}\left(Y=y|x,a,z\right)\frac{\partial\mathbb{P}\left(x|a,z\right)}{\partial x}\mathbb{P}\left(a|z\right)$ (18)
$=\sum_{a}\int dx\hskip 5.69054pt\mathbb{P}\left(Y=y|x,a\right)\frac{\partial\mathbb{P}\left(x|a,z\right)}{\partial x}\mathbb{P}\left(a|z\right)$ (19)
@@ Appendix B MIFPO and Optimal Transport
and this is independent of the variables rr! Therefore we can write the cost (14) as
@@
In this Section we prove Theorem 3.1.
To keep the notation and the main argument concise, we prove the result under the assumption that $(X,A)$ is finitely supported. Since no assumptions are made on the cardinalities of the supports, the general measurable case follows by standard approximation arguments.
We now introduce the additional notation necessary for the proof. Let $S_{0},S_{1}$ be finite disjoint sets, where $S_{a}$ represents the values of $(X,A)$ when $A=a$, for $a\in\left\{0,1\right\}$. Denote $S=S_{0}\cup S_{1}$. We are assuming that there is a probability distribution $\zeta$ on $S$, and $A$ is the random variable $A=\mathbbm{1}_{\left\{s\in S_{1}\right\}}$. $X$ is defined as taking the values $s\in S$, with $\mathbb{P}\left(X=s\right)=\zeta(s)$. Further, the variable $Y$ is defined to take values in a finite set $\mathcal{Y}$, and for every $s\in S$, its conditional distribution is given by $\rho_{s}\in\Delta_{\mathcal{Y}}$. That is, $\mathbb{P}\left(Y=y|X=s,A=a\right)=\rho_{s}(y)=\rho_{s,a}(y)$.  [^2] Note that there is a slight redundancy in the notation $\mathbb{P}\left(Y=y|X=s,A=a\right)$ here, since $a$ is determined by $s$. However, to retain compatibility with the standard notation, literature, we specify them both. This is similar to the continuous situation, in which although $A$ is technically part of the features, $X$ and $A$ are specified separately. This completes the description of the data model.
We now describe the representation. The representation will take values in a finite set $\mathcal{Z}$. For every $s\in S$ and $z\in\mathcal{Z}$, let $T_{a}(z,s)=\mathbb{P}\left(Z=z|X=s,A=a\right)$ be the conditional probability of representing $s$ as $z$. $T_{a}$ are sometimes referred as the transition kernels of the representation. For fixed $(X,A)$, the $T_{a}$’s fully define the distribution of the representation $Z$ and we shall refer to the representation as $T$ or as $Z$ in interchangeably. Finally, for $a\in\left\{0,1\right\}$ denote
With the new notation, a representation $T$ is invertible if for every $z\in\mathcal{Z}$ and every $a\in\left\{0,1\right\}$, there is at most one $s\in S_{a}$ such that $T_{a}(z,s)>0$. In words, a representation is invertible, if any given $z$ can be produced by at most two original features $s$, and at most one in each of $S_{0}$ and $S_{1}$.
Proof Of Theorem 3.1.
@@
[…]
All values of $T^{\prime}$ that were not explicitly defined in (28) are set to 0. In words, on the side of $S_{0}$, we move all the parents of $z$ except $x$ to be the parents of $z^{\prime\prime}$, while $z^{\prime}$ will have a single parent, $x$. On the $S_{1}$ side, both $z^{\prime}$ and $z^{\prime\prime}$ will have the same parents as $z$, with transitions multiplied by $\kappa$ and $1-\kappa$ respectively. The multiplication by $\kappa$ is crucial for showing both inequlaities in (10).
@@
{ $\begin{cases}\text{ for $a=1$}&\mathbb{P}\left(s|z,a\right)=\mathbb{P}\left(s|z^{\prime},a\right)=\mathbb{P}\left(s|z^{\prime\prime},a\right)\hskip 5.69054pt\forall s\in S\\ \text{ for $a=0$, $s\in U^{r}$}&\begin{cases}\mathbb{P}\left(x|z^{\prime},a\right)=1&\mathbb{P}\left(x|z^{\prime\prime},a\right)=0\\ \mathbb{P}\left(s|z^{\prime},a\right)=0&\mathbb{P}\left(s|z^{\prime\prime},a\right)=(1-\kappa)^{-1}\mathbb{P}\left(s|z,a\right)\end{cases}\end{cases}$ (31)
@@
$\sum_{s\in S_{1}}\mathbb{P}\left(Y=y|s,a=1\right)\mathbb{P}\left(s|a=1,z\right)=\sum_{s\in S_{1}}\mathbb{P}\left(Y=y|s,a=1\right)\mathbb{P}\left(s|a=1,z^{\prime}\right)$
$=\sum_{s\in S_{1}}\mathbb{P}\left(Y=y|s,a=1\right)\mathbb{P}\left(s|a=1,z^{\prime\prime}\right).$ (62)
@@ Appendix D Uniform Approximation and Two Point Representations
As discussed in Sections 1,4.1, we are interested in showing that all optimal invertible representations, no matter which, and no matter on which set $\mathcal{Z}^{\prime}$, can be approximated using a representation with the following property: For every $u\in S_{0},v\in S_{1}$, there are at most $k$ points $z\in\mathcal{Z}$ that have $(u,v)$ as parents, see Figure 2(a). Here $k$ would depend only on the desired approximation degree, but not on $\mathcal{Z}^{\prime}$, or on the exact representation we are approximating. We therefore refer to this result as the Uniform Approximation result. Its implications for practical use were discussed in Section 4.1.
The notation used in this Section was introduced in the beginning of Section C.
@@
This Section uses the notation of Section 3.3. Let $T$ be an invertible representation, let $u\in S_{0},v\in S_{1}$ be some points, and denote by $\mathcal{Z}_{uv}=\left\{z^{j}\right\}_{1}^{k}$ the set of all points $z\in\mathcal{Z}$ which have $u$ and $v$ as parents. Denote by
@@
the total weights of $\beta_{0}$ and $\beta_{1}$ transferred by the representation from $u$ and $v$ respectively to $\mathcal{Z}_{uv}$. Recall that $\rho_{u},\rho_{v}$ denote the distributions of $Y$ conditioned on $u,v$. We call the situation above, i.e. the collection of numbers $\left(\left\{\beta_{0}(u)T_{0}(z^{j},u)\right\}_{j\leq k},\left\{\beta_{1}(v)T_{1}(Z^{j},v)\right\}_{j\leq k}\right)$, a two point representation, since it describes how the weight from the points $u,v$ is distributed in the representation, independently of the rest of the representation. The contribution of $\mathcal{Z}_{uv}$ to the global performance cost is
@@
For every $\varepsilon>0$, there a number $n=n_{\varepsilon}$ depending only on the function $h$, with the following property: For every two-point representation $\left(\left\{\beta_{0}(u)T_{0}(z^{j},u)\right\}_{j\leq k},\left\{\beta_{1}(v)T_{1}(Z^{j},v)\right\}_{j\leq k}\right)$, with total weights $w_{u},w_{v}$, there is a two point representation $T^{\prime}$ on a set $\mathcal{Z}^{\prime}_{u,v}$, with the same total weights, such that $\left|\mathcal{Z}^{\prime}_{uv}\right|\leq n$, and such that
@@ Appendix E Concavity Of $E_{r}$
Appendix F MIFPO Equality Constraint
As noted in the main text, although the inequality constraint (16) is convex in the variables $r$, and can be incorporated directly into most optimisation frameworks, it may be significantly more convenient to work with equality constraints. Using the following Lemma, we can find equivalent equality constraints in a particularly simple form.
Lemma F.1.
The proof may be found in Section G.
As consequence of this result, if we find distributions $\phi_{0},\phi_{1}\in\Delta_{\mathcal{Z}}$ such that $\mu_{0}+\gamma\phi_{0}=\mu_{1}+\gamma\phi_{1}$ holds, then we know that (16) also holds, and conversely, if (16) holds, then distributions as above exist.
Using this observation, we introduce new variables, $\phi_{u,v,j}^{0}$ and $\phi_{u,v,j}^{1}$, for every $(u,v,j)\in\mathcal{Z}$, which correspond to $\phi_{0}((u,v,j))$ and $\phi_{1}((u,v,j))$ respectively. These variables will be required to satisfy the following constraints:
$\phi_{u,v,j}^{0}\geq 0,\hskip 5.69054pt\phi_{u,v,j}^{1}\geq 0\hskip 5.69054pt\forall(u,v,j)\in\mathcal{Z}$ (97)
[…]
$\beta_{0}(u)r_{u,v,j}^{0}$ $+\gamma\phi_{u,v,j}^{0}=\beta_{1}(v)r_{v,u,j}^{0}+\gamma\phi_{u,v,j}^{1}\hskip 5.69054pt\forall(u,v,j)\in\mathcal{Z}.$ (99)
Appendix G Additional Proofs
Proof Of Lemma F.1
@@ Appendix G Additional Proofs
Appendix H Experiments
This section describes additional evaluation details and experiments with fair classifiers. In SectionH.1 we provide the main algorithm figure, and discuss technical implementation details. Section H.2 contains the comparison to a number of fair classifiers, and Section H.3 discusses implementation of the entropy cost $h$ within the DCCP framework.
H.1 Implementations and computational details
Algorithm 1 MIFPO Implementation
2. Construct the histograms $\left\{\beta_{a}(l)\right\}_{l=1}^{L}$,  $a\in\left\{0,1\right\}$, for the sets  $H_{a}=\left\{c_{a}(x)\hskip 5.69054pt|\hskip 5.69054pt(x,a)\in D_{a}\right\}\subset[0,1]$.  Choose bin representatives $\left\{\rho_{l}\right\}_{l=1}^{L}$
3. Solve MIFPO, given by Definition 4.1,  with parameters $k$ and  $\left\{\beta_{a}(l)\right\}_{l=1}^{L}$, $\left\{\rho_{l}\right\}_{l=1}^{L}$, $\alpha_{a}=\left|\left\{i\hskip 5.69054pt|\hskip 5.69054pta_{i}=a\right\}\right|/N$.
The Pareto front evaluation requires two main parts - building a calibrated classifier required for evaluating $c_{a}=P(Y|X,A=a)$, and later solving the optimization problem MIFPO (see Algorithm 1).
H.2 Additional Evaluations
Following the equivalence between the Pareto fronts of fair binary classification and representations for the accuracy cost (Section 3.1), we evaluated the accuracy-fairness trade-off for some common fair classification algorithms. For our evaluation, we selected the most widely used algorithms based on GitHub repository stars and citation counts in the literature, demonstrating the importance of our proposed method in comparison to common approaches. We also evaluate FairFront, a more recent approach introduced in Wang et al., (2023).
Fair classifiers generally fall into pre-processing, in-processing, and post-processing categories. Pre- and post-processing types often utilize standard classifiers as part of their fair classification pipeline. We specifically evaluated two widely adopted algorithms, representing different categories: FairGBM Cruz et al., (2023): An in-processing method where a boosting trees algorithm (LightGBM) is subjected to pre-defined fairness constraints. Balanced-Group-Threshold Jang et al., (2021): A post-processing method which adjusts the threshold per group to obtain a fairness criterion. For FairGBM, we used the original implementation provided by the authors. For Balanced-Group-Threshold post-processing, we utilized implementations available via Aequitas Saleiro et al., (2018), a popular bias and fairness audit toolkit.
We conducted our evaluation using three of the most common datasets in this field, which are known to have relevance to real-world decision-making processes: the Adult dataset (income prediction), COMPAS (recidivism prediction), and LSAC (law school admission).
It is important to note that, as a rule, common fairness classification methods are not designed to control the fairness-accuracy trade-off explicitly. Instead, in most cases, these methods rely on rerunning the algorithm for a wide range of hyperparameter settings, in the hope that different hyperparameters would result in different fairness-accuracy trade-off points. However, there typically is no direct known and controlled relation between hyperpatameters and the obtained fairness-accuracy trade-off. For FairGBM, we utilized the hyperparameter ranges specified in the original paper, Cruz et al., (2023). In the case of the balancing post-processing method, we conducted a grid search over the full range of all possible hyperparameters to ensure a comprehensive analysis.
Figure 4 shows the MIFPO computed Pareto front, and all hyperparameter runs of the two algorithms above, with accuracy evaluated on the test set. These experiments demonstrate the following two points: (a) The standard classifiers achieve a considerably lower accuracy than what is theoretically possible at a given fairness level. (b) the existing methods are also unable to present solutions for the full range of the statistical parity values. The values from the FGBM and the post-processing algorithms all have statistical parity $\leq 0.2$. Similarly to to the case of fair representations, these results emphasize the limitations of current fair classifiers in achieving optimal trade-offs between accuracy and fairness across the full range of fairness values.
Figure 5: Comparing FairFront to MIFPO accuracy-fairness tradeoff on two curated datasets.
Additionally, we add a comparison to FairFront (Wang et al.,, 2024), which, similarly to our work, depends on estimates of $P(Y|X)$. See the discussion in Section 2. We note, however, that the utility of this comparison is limited due to the very strict setup defined in that paper, which cannot be applied to natural datasets. Specifically: 1) The implementation in Wang et al., (2023) requires creating a finite discrete space by binning over the full $\mathbb{R}^{d}$ feature space [^3] This is completely different from the discretizaton of $\Delta_{\mathcal{Y}}$ used in MIFPO. The space $\mathbb{R}^{d}$ is the feature space, and is much larger than $\Delta_{\mathcal{Y}}$., which allows for perfect modeling of the distribution $P(Y|X)$ (by simple counting). This is not normally possible on real datasets in practice. 2) The binning itself was performed manually for each dataset separately, and we were unable to discern the logic behind the selected parameters. 3) To allow reasonable binnig, the true dimensions of the data are manually reduced, again by picking features manually for each dataset. Due to these issues, it was practically impossible to apply the method in Wang et al., (2023) to other standard datasets, or even to the datasets used in Wang et al., (2023) but with standard features.
Nonetheless, for the sake of comparison, we compared MIFPO and FairFront on the same restricted, preprocessed version of the datasets that FairFront used, with the same bin-based probability estimator. The results are shown in Figure 5.
@@ H.2 Additional Evaluations
H.3 Minimization of Concave Functions under Convex Constraints and the Entropy Loss
@@
We can write the cost function under the entropy accuracy error as:
@@
$\mathrm{entropy}(a+b-a\cdot p_{a}-b\cdot p_{b})+\mathrm{entropy}(a\cdot p_{a}+b\cdot p_{b})-\mathrm{entropy}(a+b)$
$\mathrm{entropy}(a\cdot(1-p_{a})+b\cdot(1-p_{b}))+\mathrm{entropy}(a\cdot p_{a}+b\cdot p_{b})-\mathrm{entropy}(a+b)$
@@
Appendix I Fair Classifiers As Fair Representations
As discussed in Section 3.1, Pareto front of binary classifiers with statistical parity can be computed from the Pareto front of representations with total variation fairness distance. In this Section we provide the proof of this result, Lemma 3.1. The Lemma and the related notation are restated here for convenience.